# Towards Multi-Modal Animal Pose Estimation: An In-Depth Analysis

### v1

@@ 6. Multi-Modal Animal Pose Estimation
Researchers have begun to explore the integration of multi-sensor and multi-modality to improve the accuracy, robustness, and ecological validity of APE algorithms (Patel et al., 2017; Kearney et al., 2020a; Li et al., 2022; Luo et al., 2023; Patel et al., 2023; Zhang et al., 2023; Ye et al., 2023; Muramatsu et al., 2024; Deng, 2023). In this section, we examine the motivation behind multi-modal systems, review existing multi-modal HPE and APE techniques, discuss how they tackle the limitations in unimodal APE methods, and highlight applications demonstrating the potential of integrating diverse sensors and modalities for APE, as shown in table 8.
6.1. Multi-Modal System in Human Pose Estimation
@@
To inspire the use of multi-modal systems for APE, we draw insights from recent multi-modal HPE studies. For instance, (Zheng et al., 2022; Zanfir et al., 2023) combine data from visual RGB cameras with depth information obtained from RGBD and LiDAR sensors in autonomous driving to increase the robustness of 3D HPE in complex environments. The mRI dataset (An et al., 2022) integrates mmWave radar, RGBD cameras, and IMUs and provides a benchmark to support 3D HPE sensor fusion research. Another study (An et al., 2022) introduces the mRI dataset, a multi-modal dataset for 3D HPE that integrates data from mmWave radar, RGBD cameras, and IMUs. These studies highlight the benefits of multi-modal integration in HPE and further inspire us to explore similar approaches for APE.
6.2. Motivation for Multi-Modal System in Animal Pose Estimation
@@
Unimodal APE faces challenges like occlusions and limited perspectives, which multi-modal systems can overcome by combining complementary data to improve accuracy and understanding of animal behaviour (Deng, 2023). For example, multi-modal systems enhance APE resilience by compensating for sensor failures, such as using thermal in low-light conditions when RGB cameras struggle. In addition, RGB-based cameras excel in capturing detailed visual information but may falter with occlusions, while some mmWave sensors can see through walls, and LiDAR can provide valuable 3D spatial data; these significantly help us to expand the coverage area and reduce blind spots. Beyond visual data, integrating IMUs and audio sensors further enriches animal behaviour analysis by capturing motion and auditory cues, which makes multi-modal systems essential for a more robust APE and a better understanding of dynamic environments.
6.3. Existing Multi-Modal Animal Pose Estimation Techniques
@@
In recent years, there has been growing interest in the development of multi-modal APE systems (Kearney et al., 2020a; Luo et al., 2023; Patel et al., 2017; Li et al., 2022; Patel et al., 2023; Zhang et al., 2023; Ye et al., 2023; Muramatsu et al., 2024; Li et al., 2024a; Zuffi and Black, 2024). It involves combining information from different sensors to derive a unified representation of animal pose. Compared to the significant progress in unimodal APE, there is limited research on using multiple sensors and modalities for enhanced accuracy in APE, and we aim to review all of the existing multi-modal APE studies in this section.
Animal-borne cameras + GPS + IMUs (Ego-view)
@@
One notable example of a multi-modal system using animal-borne cameras, GPS, and IMU is the work on tracking the cheetah tail (Patel et al., 2017) in ego-view. Ego-view 3D APE, though less explored, presents a novel perspective by placing cameras on the animals themselves, offering a first-person viewpoint. In this study (Patel et al., 2017), researchers attached camera and sensor equipment to cheetahs to obtain motion data from the animal’s viewpoint. Using kinematic models, they can estimate the 3D pose of the cheetah’s tail, including its position, orientation, and movement in 3D space. This research demonstrates the potential of multi-modal ego-view APE methods in capturing the animal’s pose and movement directly from the animal’s perspective, which helps us gain insights into the animal’s behaviour, locomotion, and agility so we can have a better understanding of the cheetah’s hunting strategies, balance control, and communication signals. Please note that the ego-view 3D APE study is currently limited, and the cheetah tail tracking study opens a door for it.
RGB + Depth
@@
One promising sensor modality is the RGBD-based pose estimation, as exemplified by (Kearney et al., 2020a). The pose estimation model is a two-stage approach for predicting 3D canine poses from RGBD data. In the first stage, a stacked hourglass network is employed to predict 2D joint locations from RGB images, with multiple stacked layers that allow for iterative refinement of joint positions. To generalise the model across various canine shapes and sizes, this study trains the network on both real and synthetic data. In the second stage, a Hierarchical Gaussian Process Latent Variable Model (H-GPLVM) is used to lift the predicted 2D joint positions to 3D, while ensuring the output poses are anatomically plausible by imposing constraints on the possible configurations. The H-GPLVM leverages the structure of the canine skeleton to refine the pose estimations, which further reduces errors and improves robustness in complex poses or challenging viewpoints. This model can extract dog poses effectively even when there is significant variability in appearance and motion. In 2023, there is another RGBD-based APE approach (Luo et al., 2023) that enhances livestock body measurements. It first preprocesses point cloud data to detect 3D keypoints, then aligns a template mesh (i.e. a statistical shape model) with the scan data, and uses non-rigid registration and poses normalisation to standardise the animals’ poses into a canonical form. These help to address the common challenges such as noise, missing data, and inconsistencies caused by livestock movement. Even though RGBD is a promising approach, its popularity is limited by dataset availability and hardware constraints; hence, it has not yet been widely adopted within the APE community compared to RGB-based methods.
RGB + Depth + Thermal infrared
@@
The OptiPose study (Patel et al., 2023) introduces a novel approach for animal pose tracking that leverages a multi-modal dataset and token-based pose optimisation. To capture accurate APE, researchers collected a 3D multi-modal dataset comprising RGB, depth, and thermal infrared data. RGB data provides high-resolution colour imagery that helps identify detailed animal features, allowing for better 2D keypoint extraction. Depth data provides spatial information about the distance between the camera and the animal, which is important for reconstructing the 3D pose from multiple 2D views. Finally, thermal infrared data measures the heat emitted by the animal, which is particularly useful in low-light or occluded conditions where RGB data struggled. They propose a token-based optimisation method that uses multi-modal data to refine the pose estimation results. The study demonstrates the potential of integrating multiple sensor modalities for improved animal pose tracking and motivates further exploration of multi-modal approaches in animal behaviour analysis.
RGB + LiDAR
@@
A novel approach to APE uses LiDAR and RGB sensors, as exemplified by the WildPose (Muramatsu et al., 2024) system. Unlike traditional methods that rely on RGB images, WildPose utilises a combination of a zoom-lens camera and a solid-state LiDAR to capture both 2D and 3D data of animals in their natural habitats. The WildPose system is non-invasive and can capture creatures in a long-range distance (ranging from 2-200 meters) without disturbing the animals. Moreover, LiDAR adds depth information about the animals and their surrounding environments to provide a more comprehensive understanding of the animal’s morphology and movements. Despite these advantages, the high cost and complexity of LiDAR systems may limit their accessibility and scalability. Additionally, the accuracy and quality of the LiDAR data can be affected by environmental factors such as weather conditions and lighting.
RGB + Acoustic
@@
The Sound of Motion (Li et al., 2022) and CLHOP (Li et al., 2024a) present a multi-modal approach for estimating horse motion using video and audio data. Audio features are extracted from sounds like horse hoof strikes and are processed in parallel with visual features. These audio features are converted into Mel spectrograms and passed through a network that helps predict 3D pose parameters. These studies highlight the value of incorporating audio to capture fine-grained details of animal movements. Apart from APE, audiovisual data have also been applied in other animal motion analyses such as horse behaviour detection (Nunes et al., 2021) and primate action recognition (Bain et al., 2021).
RGB + Language
@@
As we advance, the APE trend is moving towards using vision-language-based APE methods (Zhang et al., 2023; Zuffi and Black, 2024). The CLAMP (2023) (Zhang et al., 2023) method uses language to enhance learning processes through the description provided, thus increasing the capability of generalising models about new and unseen animal poses. Following that, the AWOL (2024) (Zuffi and Black, 2024) approach creates a model that takes input from Contrastive Language-Image Pretraining (CLIP) encodings of either text or images and maps them to 3D shape parameters. This learns the mapping between the CLIP latent space and the shape space of SMAL, allowing it to generate realistic 3D animal mesh as a final output. This exploits the strengths of vision-language models in blending visual and textual information into a more robust framework, advancing the APE field of work.
@@
7. Future Work
Can we reduce or eliminate the effort required for ground-truth annotation in APE using unsupervised learning and meta-learning?
@@ 7. Future Work
How can we extend multi-modal frameworks to perform multiple-animal pose estimation in the scene?
@@
Currently, the use of multi-modal frameworks for a single animal is limited, and for multi-animal, minimal work exists (i.e. OptiPose (Patel et al., 2023) that can perform APE on two rodents from a multi-modal system). However, the future work will include multi-modal multi-animals pose estimation in-the-wild for various species. Going forward, we will see a fusion of sensors and modalities for handling occlusions and provide additional context when animals interact closely. For example, depth data will help to differentiate between overlapping animals, while thermal data will assist in distinguishing animals from their surroundings in low-light conditions. In addition, we may also use multi-modal data to combine animal appearance and other information such as their shape, weight, and gaits together to resolve individual animal identity, which is very important for agriculture and conservation. In future, work will move towards scene object interaction and reconstruction. By using multi-modal data, we will reconstruct both the animals and their surrounding objects by understanding context-specific interactions, such as animals navigating around obstacles, interacting with other animals, or using objects (e.g. chimpanzees nut cracking in-the-wild). For instance, LiDAR or depth sensors will capture the 3D geometry of the environment, and when combined with RGB data, the system will reconstruct both the animal’s poses and the objects they interact with. The future will also move towards having real-time processing, which will allow us to track pose and behaviour in real-time, enabling effective health monitoring to protect endangered species.
@@
Can we perform language-driven animal articulation with advanced APE methods?
@@
With the rise of models like CLIP, the trend is moving towards more vision-language hybrid approaches within the CV community. So far, there is no work available where it is possible to control and articulate 3D animal models using natural language commands in an unsupervised manner, but this will likely happen in the future. By integrating language-driven models with advanced APE techniques, we can achieve sophisticated text-based control for animal poses. Similar to how text prompts can control image generation in models like CLIP, future APE methods could allow users to articulate an animal’s pose using natural language instructions. For instance, commands like ”raise the horse’s front right leg” or ”make the chimpanzee sit” could be translated into precise joint movements or skeletal adjustments in a 3D model. Moreover, mesh-based APE methods, when integrated with language models, could allow for real-time adjustments in the animal’s anatomy and surface details with commands like ”increase muscle tension” or ”broaden the horse’s stance” and this could instantly alter both the pose and physical appearance of the 3D mesh model in real-time. These advancements will be beneficial in applications like virtual reality or animations, where controlling animal movements dynamically with text could simplify workflows. It also has the potential for behavioural simulations; for example, researchers or animators can generate lifelike animal behaviours through high-level language inputs without manually defining joint movements.
Acknowledgements.
@@
We thank Andrew Zisserman for his excellent suggestions and for reviewing the paper. We would also like to thank Jiaxing Zhong, Ta-Ying Cheng, Xinyu Hou, Qian Xie, Shaokai Ye, Daniel Schofield, Eldar Insafutdinov, Tomas Jakab and Anjun Hu for the insightful discussion on pose estimation. We would like to thank (Yunus et al., 2024) for providing the symbols. Additionally, Qianyi would like to thank Jian Zhou for his loving support. During this project, Qianyi Deng was supported by the UKRI’s EPSRC ACE-OPS grant (EP/S030832/1), and Oishi Deb was supported by the UKRI’s EPSRC AIMS CDT grant (EP/S024050/1).

### v2

@@ 6 Multi-Modal Animal Pose Estimation
Researchers have begun to explore the integration of multi-sensor and multi-modality to improve the accuracy, robustness, and ecological validity of APE algorithms [108, 56, 67, 84, 109, 169, 161, 97, 24]. In this section, we examine the motivation behind multi-modal systems, review existing multi-modal HPE and APE techniques, discuss how they tackle the limitations in unimodal APE methods, and highlight applications demonstrating the potential of integrating diverse sensors and modalities for APE, as shown in Table 8.
6.1 Multi-Modal System in Human Pose Estimation
@@
To inspire the use of multi-modal systems for APE, we draw insights from recent multi-modal HPE studies. For instance, [172, 167] combine data from visual RGB cameras with depth information obtained from RGBD and LiDAR sensors in autonomous driving to increase the robustness of 3D HPE in complex environments. The mRI dataset [3] integrates mmWave radar, RGBD cameras, and IMUs and provides a benchmark to support 3D HPE sensor fusion research. Another study [3] introduces the mRI dataset, a multi-modal dataset for 3D HPE that integrates data from mmWave radar, RGBD cameras, and IMUs. These studies highlight the benefits of multi-modal integration in HPE and further inspire us to explore similar approaches for APE.
6.2 Motivation for Multi-Modal System in Animal Pose Estimation
@@
Unimodal APE faces challenges like occlusions and limited perspectives, which multi-modal systems can overcome by combining complementary data to improve accuracy and understanding of animal behaviour [24]. For example, multi-modal systems enhance APE resilience by compensating for sensor failures, such as using thermal in low-light conditions when RGB cameras struggle. In addition, RGB-based cameras excel in capturing detailed visual information but may falter with occlusions, while some mmWave sensors can see through walls, and LiDAR can provide valuable 3D spatial data; these significantly help us to expand the coverage area and reduce blind spots. Beyond visual data, integrating IMUs and audio sensors further enriches animal behaviour analysis by capturing motion and auditory cues, which makes multi-modal systems essential for a more robust APE and a better understanding of dynamic environments.
6.3 Existing Multi-Modal Animal Pose Estimation Techniques
@@
In recent years, there has been growing interest in the development of multi-modal APE systems [56, 84, 108, 67, 109, 169, 161, 97, 66, 175]. It involves combining information from different sensors to derive a unified representation of animal pose. Compared to the significant progress in unimodal APE, there is limited research on using multiple sensors and modalities for enhanced accuracy in APE, and we aim to review all of the existing multi-modal APE studies in this section.
@@
Animal-borne cameras + GPS + IMUs (Ego-view): One notable example of a multi-modal system using animal-borne cameras, GPS, and IMU is the work on tracking the cheetah tail [108] in ego-view. Ego-view 3D APE, though less explored, presents a novel perspective by placing cameras on the animals themselves, offering a first-person viewpoint. In this study [108], researchers attached camera and sensor equipment to cheetahs to obtain motion data from the animal’s viewpoint. Using kinematic models, they can estimate the 3D pose of the cheetah’s tail, including its position, orientation, and movement in 3D space. This research demonstrates the potential of multi-modal ego-view APE methods in capturing the animal’s pose and movement directly from the animal’s perspective, which helps us gain insights into the animal’s behaviour, locomotion, and agility so we can have a better understanding of the cheetah’s hunting strategies, balance control, and communication signals. Please note that the ego-view 3D APE study is currently limited, and the cheetah tail tracking study opens a door for it.
@@
RGB + Depth: One promising sensor modality is the RGBD-based pose estimation, as exemplified by [56]. The pose estimation model is a two-stage approach for predicting 3D canine poses from RGBD data. In the first stage, a stacked hourglass network is employed to predict 2D joint locations from RGB images, with multiple stacked layers that allow for iterative refinement of joint positions. To generalise the model across various canine shapes and sizes, this study trains the network on both real and synthetic data. In the second stage, a Hierarchical Gaussian Process Latent Variable Model (H-GPLVM) is used to lift the predicted 2D joint positions to 3D, while ensuring the output poses are anatomically plausible by imposing constraints on the possible configurations. The H-GPLVM leverages the structure of the canine skeleton to refine the pose estimations, which further reduces errors and improves robustness in complex poses or challenging viewpoints. This model can extract dog poses effectively even when there is significant variability in appearance and motion. In 2023, there is another RGBD-based APE approach [84] that enhances livestock body measurements. It first preprocesses point cloud data to detect 3D keypoints, then aligns a template mesh (i.e. a statistical shape model) with the scan data, and uses non-rigid registration and poses normalisation to standardise the animals’ poses into a canonical form. These help to address the common challenges such as noise, missing data, and inconsistencies caused by livestock movement. Even though RGBD is a promising approach, its popularity is limited by dataset availability and hardware constraints; hence, it has not yet been widely adopted within the APE community compared to RGB-based methods.
@@
RGB + Depth + Thermal infrared: The OptiPose study [109] introduces a novel approach for animal pose tracking that leverages a multi-modal dataset and token-based pose optimisation. To capture accurate APE, researchers collected a 3D multi-modal dataset comprising RGB, depth, and thermal infrared data. RGB data provides high-resolution colour imagery that helps identify detailed animal features, allowing for better 2D keypoint extraction. Depth data provides spatial information about the distance between the camera and the animal, which is important for reconstructing the 3D pose from multiple 2D views. Finally, thermal infrared data measures the heat emitted by the animal, which is particularly useful in low-light or occluded conditions where RGB data struggled. They propose a token-based optimisation method that uses multi-modal data to refine the pose estimation results. The study demonstrates the potential of integrating multiple sensor modalities for improved animal pose tracking and motivates further exploration of multi-modal approaches in animal behaviour analysis.
@@
RGB + LiDAR: A novel approach to APE uses LiDAR and RGB sensors, as exemplified by the WildPose [97] system. Unlike traditional methods that rely on RGB images, WildPose utilises a combination of a zoom-lens camera and a solid-state LiDAR to capture both 2D and 3D data of animals in their natural habitats. The WildPose system is non-invasive and can capture creatures in a long-range distance (ranging from 2-200 meters) without disturbing the animals. Moreover, LiDAR adds depth information about the animals and their surrounding environments to provide a more comprehensive understanding of the animal’s morphology and movements. Despite these advantages, the high cost and complexity of LiDAR systems may limit their accessibility and scalability. Additionally, the accuracy and quality of the LiDAR data can be affected by environmental factors such as weather conditions and lighting.
@@
RGB + Acoustic: The Sound of Motion [67] and CLHOP [66] present a multi-modal approach for estimating horse motion using video and audio data. Audio features are extracted from sounds like horse hoof strikes and are processed in parallel with visual features. These audio features are converted into Mel spectrograms and passed through a network that helps predict 3D pose parameters. These studies highlight the value of incorporating audio to capture fine-grained details of animal movements. Apart from APE, audiovisual data have also been applied in other animal motion analyses such as horse behaviour detection [104] and primate action recognition [5].
@@
RGB + Language: As we advance, the APE trend is moving towards using vision-language-based APE methods [169, 175]. The CLAMP (2023) [169] method uses language to enhance learning processes through the description provided, thus increasing the capability of generalising models about new and unseen animal poses. Following that, the AWOL (2024) [175] approach creates a model that takes input from Contrastive Language-Image Pretraining (CLIP) encodings of either text or images and maps them to 3D shape parameters. This learns the mapping between the CLIP latent space and the shape space of SMAL, allowing it to generate realistic 3D animal mesh as a final output. This exploits the strengths of vision-language models in blending visual and textual information into a more robust framework, advancing the APE field of work.
@@
7 Future Work
7.1 Can we reduce or eliminate the effort required for ground-truth annotation in APE using unsupervised learning and meta-learning?
@@ 7 Future Work
7.2 How can we extend multi-modal frameworks to perform multiple-animal pose estimation in the scene?
@@
Currently, the use of multi-modal frameworks for a single animal is limited, and for multi-animal, minimal work exists (i.e. OptiPose [109] that can perform APE on two rodents from a multi-modal system). However, the future work will include multi-modal multi-animals pose estimation in-the-wild for various species. Going forward, we will see a fusion of sensors and modalities for handling occlusions and provide additional context when animals interact closely. For example, depth data will help to differentiate between overlapping animals, while thermal data will assist in distinguishing animals from their surroundings in low-light conditions. In addition, we may also use multi-modal data to combine animal appearance and other information such as their shape, weight, and gaits together to resolve individual animal identity, which is very important for agriculture and conservation. In future, work will move towards scene object interaction and reconstruction. By using multi-modal data, we will reconstruct both the animals and their surrounding objects by understanding context-specific interactions, such as animals navigating around obstacles, interacting with other animals, or using objects (e.g. chimpanzees nut cracking in-the-wild). For instance, LiDAR or depth sensors will capture the 3D geometry of the environment, and when combined with RGB data, the system will reconstruct both the animal’s poses and the objects they interact with. The future will also move towards having real-time processing, which will allow us to track pose and behaviour in real-time, enabling effective health monitoring to protect endangered species.
@@
7.3 Can we perform language-driven animal articulation with advanced APE methods?
@@
With the rise of models like CLIP, the trend is moving towards more vision-language hybrid approaches within the CV community. So far, there is no work available where it is possible to control and articulate 3D animal models using natural language commands in an unsupervised manner, but this will likely happen in the future. By integrating language-driven models with advanced APE techniques, we can achieve sophisticated text-based control for animal poses. Similar to how text prompts can control image generation in models like CLIP, future APE methods could allow users to articulate an animal’s pose using natural language instructions. For instance, commands like “raise the horse’s front right leg” or “make the chimpanzee sit” could be translated into precise joint movements or skeletal adjustments in a 3D model. Moreover, mesh-based APE methods, when integrated with language models, could allow for real-time adjustments in the animal’s anatomy and surface details with commands like “increase muscle tension” or “broaden the horse’s stance” and this could instantly alter both the pose and physical appearance of the 3D mesh model in real-time. These advancements will be beneficial in applications like virtual reality or animations, where controlling animal movements dynamically with text could simplify workflows. It also has the potential for behavioural simulations; for example, researchers or animators can generate lifelike animal behaviours through high-level language inputs without manually defining joint movements.
Declarations
@@
We thank Professor Andrew Zisserman for his excellent suggestions and for reviewing the paper. We would also like to thank Jiaxing Zhong, Ta-Ying Cheng, Xinyu Hou, Qian Xie, Shaokai Ye, Daniel Schofield, Eldar Insafutdinov, Niki Amini-Naieni, Tomas Jakab and Anjun Hu for insightful technical discussion. We would like to thank [166, 20] for using some of their symbols, templates and scripts. Additionally, Qianyi would like to thank Jian Zhou for his loving support. During this project, Qianyi Deng was supported by the UKRI’s EPSRC ACE-OPS grant (EP/S030832/1), and Oishi Deb was supported by the UKRI’s EPSRC AIMS CDT grant (EP/S024050/1).